How many grams of dry NH4Cl need to be added to 1.70 L of a 0.500 M solution of ammonia, NH3, to prepare a buffer solution that has a pH of 8.80? Kb for ammonia is 1.8×10−5.
Answer #1

moles of NH3 = 1.70 x 0.500 = 0.85

pH = 8.80

pOH = 14 - 8.80 = 5.20

Kb = 1.8 x 10^-5

pKb = -log (1.8 x 10^-5)

        = 4.74

pOH = pKb + log [salt / base]

5.20 = 4.74 + log [NH4Cl / NH3]

5.20 = 4.74 + log [NH4Cl / 0.85]

moles of NH4Cl = 2.45 mol

mass of NH4Cl = 2.45 x 53.49

mass of NH4Cl   = 131 g